Run Meeting: Feb 26th, 2019
From Hall A Wiki
Back to:
3:30PM Run Meeting, Counting house meeting room (200C)
RC: Cameron Clarke
Contents
Status
Yesterday Day and Swing
- Left Dipole restored at the end of Day shift: https://logbooks.jlab.org/entry/3659482
- Looking at RPI communication issues, potential fix by Ed Brash: https://logbooks.jlab.org/entry/3659348
- Bob Michaels investigated questionable DAQ wiring and added additional path to a new scaler for T2, T5, and T6 as well as added T6 to the RHRS digital oscope: https://logbooks.jlab.org/entry/3659475
- Ongoing analysis questions about timing and whether things are ok: https://logbooks.jlab.org/entry/3659497
- Immediately after beam was restored following the Left Dipole repair work we noticed the scalers were acting crazy - specifically that T6 was fluctuating wildly, and with no beam was ~= T5, indicating a crossed wire: https://logbooks.jlab.org/entry/3659540, https://logbooks.jlab.org/entry/3659554
- Alexandre Camsonne took access and swapped out the JLab custom level translator that handles the TS and scaler trigger signal paths: https://logbooks.jlab.org/entry/3659567
- This appears to have fixed the problem of wildly fluctuating rates in scaler gui with beam on: https://logbooks.jlab.org/entry/3659585
- Though we now have noticed that for the last two days the RHRS T2 signal frequency is 2x larger than the LHRS T2, meaning double pulsing probably: https://logbooks.jlab.org/entry/3659790, https://logbooks.jlab.org/entry/3659593, https://logbooks.jlab.org/entry/3659610#comment-20417, https://logbooks.jlab.org/entry/3659693
- The new module has probably caused a timing shift of 10 ns or so in all signals going through it, under investigation if true and OK by Ranit and Bob: https://logbooks.jlab.org/entry/3659612, https://logbooks.jlab.org/entry/3659673,
- Plan to access during today's beam studies to evaluate and fix this problem
- Silviu simulates and we are eventually authorized for 2x2 raster, optics 1 target, 50 uA max limit running for future conditions: https://logbooks.jlab.org/entry/3659587
- Good 40, 45, and 50 uA running during Swing: https://logbooks.jlab.org/entry/3659722
Today Owl
- 50 uA production running: https://logbooks.jlab.org/entry/3659926
- Accidentally left the ps7=1, so the data will also have cosmics: https://logbooks.jlab.org/entry/3659957
- 5 AM transients in Helium coolant levels noticed by Ryan Richards, Jessie informed and plans to investigate at 9 AM: https://logbooks.jlab.org/entry/3659961, https://logbooks.jlab.org/entry/3659964
Today Day
- Radcon surveys the hall so that Bob and Alex can investigate RHRS T2 double pulsing question: https://logbooks.jlab.org/entry/3659970
- 50 uA production after the North Linac RF work is completed and 5 pass separator problems resolved
- BPM Pedestal run by Jason Bane: https://logbooks.jlab.org/entry/3660024
- LQ3 coolant level investigated and corrected: https://logbooks.jlab.org/entry/3659981
- Sci Fi Cosmic tests and DAQ improvement:
- Look at the DAQ
- Issues appear to be resolved: https://logbooks.jlab.org/entry/3660025,
- Added the T5 and T6 coincidences to the digital oscope channels 2 and 4: https://logbooks.jlab.org/entry/3660022
- Bad Level translator board given to Electronics Group and investigated: https://logbooks.jlab.org/entry/3660044
- Look at the Helium levels of magnets, corrected mode of PID loop
- Investigation of impact on DAQ of the time offsets and T6 extra pulsing by Bob: https://logbooks.jlab.org/entry/3660075
Hall Status
- Taking to controlled access to take advantage of RF recovery from 8AM to noon in order to investigate double pulsing in RHRS T2 signal
- Magnet coolant levels had some spikes around 5AM so investigations may need to be made during access
- Otherwise perfectly capable of accepting beam and will continue production at 50uA the rest of the day and onwards
Plan
- Continue production data on Tungsten 2.8% target at 50uA, and move higher if possible
- In the future we should avoid short notice activities in the hall outside of normal business hours, as short notice increases chances of inadequate preparation and likelihood of problems requiring hall access or calling in experts consequently increases, and the present high quality and high quantity 50 uA running is precious
RC
- In lieu of today's cancelled Hall A general meeting at 1:30, please see these plots that summarize running up until now (will be incorporated fully into Online Replay scripts before I leave RC duties tomorrow afternoon): https://logbooks.jlab.org/entry/3659616
Cameron Clarke is the RC (email cameronc@jlab.org, RC phone: 7578761787) All shifts are on-call and should expect to receive full beam for production on tungsten foils, with occasional sci-fi runs pending.
Slides/Links
MCC scheduling Whiteboard: http://accboard.acc.jlab.org/
APEX wiki page: https://hallaweb.jlab.org/wiki/index.php/APEX
Bluejeans connection information
To join the Meeting: https://bluejeans.com/256205934
To join via Room System: Video Conferencing System: bjn.vc -or-199.48.152.152 Meeting ID : 256205934
To join via phone : 1) Dial:
+1.408.740.7256 (United States) +1.888.240.2560 (US Toll Free) +1.408.317.9253 (Alternate number) (see all numbers - http://bluejeans.com/numbers)
2) Enter Conference ID : 256205934